3 out of 3 attempts Enter your answer in the provided box. Calculate the pH at the equivalence point in the titration of 30.0 mL of 0.195 M methylamine (K b 4.4 x 10 4) with 0.305 M HCl pH 4.9 As
Answer #1

1st calculate the volume of HCl required

At equivalence point,

mol of methylamine = mol of acid

0.195 M * 30.0 mL = 0.305 M * V

V = 19.2 mL

This is volume of acid added

CH3NH2 + HCl <—> CH3NH3+

mol of CH3NH3+ formed = mol of CH3NH2

= 0.195 M * 30.0 mL

=5.85 mmol

[CH3NH3+] = number of mol / total volume

= 5.85 mmol / (30.0 + 19.2) mL

= 0.119 M

Now this is weak acid

Ka = 10^-14/Kb

= 10^-14 / (4.4*10^-4)

= 2.27*10^-11

It will dissociate as,

CH3NH3+   <—>   CH3NH2   +   H+

0.119           0           0   (initial)

0.119-x       x           x   (at equilibrium)

Ka = x*x/(0.119-x)

2.27*10^-11 = x*x/(0.119-x)

since Ka is small, x will be small and it be ignored as compared to 0.119

above expression thus becomes,

2.27*10^-11 = x*x/(0.119)

x = 1.64*10^-6 M

So,

[H+] = x = 1.64*10^-6 M

pH = -log [H+]

= -log (1.64*10^-6)

= 5.78

Answer: 5.78